# Binary Cumulative Encoding meets Time Series Forecasting

### v2

@@ 5.2 Results
We evaluated the performance of five different models: BinConv (our model), GRU-NVP, TimeGrad, and PatchTST, which are briefly described in Section 2. We also added simple linear model DLinear [zeng2023transformers] for the comparison. For all baseline models, we use the optimal set of hyperparameters and normalization strategies provided by [zhang2024probts]. For BinConv, as discussed in Section 3, we used mean scaling before quantization. Mean scaling was applied independently to each sample, as the M4 datasets consist of diverse time series with varying scales. We did not fine-tune the BinConv model’s hyperparameters for each dataset, as it already demonstrated strong performance with the default configuration. The hyperparameters used across all datasets are summarized in Table 2.
Table 2: Default hyperparameters used for BinConv across all datasets. $C$ denotes the context length.
Hyperparameter Value
Number of bins ($D$) $1000$
Minimum bin value ($b_{0}$) $-5$
Maximum bin value ($b_{D}$) $5$
Number of output channels ($K$) $C$
Kernel size of 2D convolution in BinConv block ($s_{1}$) $3$
Kernel size of 1D convolution in BinConv block ($s_{2}$) $3$
Kernel size of final 1D convolution ($s_{3}$) $51$
Number of BinConv blocks ($M$) $3$
Dropout $0.35$
Learning rate $0.001$
The comparison between models is presented in Table 3. We trained and evaluated all models using five different random seeds and report the minimum, average, and maximum scores for each model. The metrics of BinConv are consistently better than those of the baseline models. Notably, the average scores of BinConv for both CRPS and NMAE are better than the best scores of all baselines.
Additionally, Table 4 compares the performance of BinConv when using argmax-based forecasting versus sampling-based forecasting (see Section 4.2 for details). While using point forecasting slightly degrades performance, it still outperforms all baseline models. We provide additional results on multivarite time-series datasets in Appendix C.
Table 3: CRPS and NMAE results for all models. Each metric is reported as the average (AVG), minimum (MIN), and maximum (MAX) over 5 different seeds.
Dataset Stat DLinear GRU-NVP PatchTST TimeGrad BinConv
@@
m4_daily AVG 0.0440 0.0440 0.0714 0.0782 0.0438 0.0438 0.0587 0.0696 0.0327 0.0382
m4_weekly AVG 0.1046 0.1046 0.0972 0.1062 0.1039 0.1039 0.1046 0.1283 0.0902 0.0972
tourism_monthly AVG 0.2745 0.2745 0.3723 0.4536 0.2633 0.2633 0.2295 0.2968 0.1824 0.1955
m4_daily MIN 0.0394 0.0394 0.0570 0.0649 0.0404 0.0404 0.0402 0.0475 0.0291 0.0341
m4_weekly MIN 0.1007 0.1007 0.0916 0.1034 0.1020 0.1020 0.0800 0.0985 0.0851 0.0921
tourism_monthly MIN 0.2595 0.2595 0.2880 0.3849 0.2281 0.2281 0.1993 0.2596 0.1777 0.1911
m4_daily MAX 0.0506 0.0506 0.1196 0.1259 0.0476 0.0476 0.0933 0.1077 0.0366 0.0427
m4_weekly MAX 0.1072 0.1072 0.1006 0.1085 0.1078 0.1078 0.1718 0.2121 0.0981 0.1055
tourism_monthly MAX 0.3077 0.3077 0.5799 0.5399 0.2888 0.2888 0.2500 0.3186 0.1892 0.2012
Table 4: Comparison of BinConv performance using sampling versus argmax forecasting. Each metric is reported as the average (AVG), minimum (MIN), and maximum (MAX) over 5 different seeds.
@@
m4_daily AVG 0.0371 0.0371 0.0327 0.0382
m4_weekly AVG 0.0975 0.0975 0.0902 0.0972
tourism_monthly AVG 0.1961 0.1961 0.1824 0.1955
m4_daily MIN 0.0348 0.0348 0.0291 0.0341
m4_weekly MIN 0.0927 0.0927 0.0851 0.0921
tourism_monthly MIN 0.1916 0.1916 0.1777 0.1911
m4_daily MAX 0.0387 0.0387 0.0366 0.0427
m4_weekly MAX 0.1049 0.1049 0.0981 0.1055
tourism_monthly MAX 0.2013 0.2013 0.1892 0.2012
@@ 5.3 Efficiency of the Model
m4_daily $1.204$ $20.173$ $47.470$ $107.434$ $141.108$
m4_weekly $1.040$ $17.680$ $43.725$ $107.434$ $141.108$
tourism_monthly $3.504$ $54.013$ $155.288$ $107.434$ $141.108$
@@
Train $\mathbf{0.0123\pm 0.0026}$ $0.0284\pm 0.0043$ $0.0284\pm 0.0043$ $0.0381\pm 0.0049$ $0.0450\pm 0.0063$ $0.0362\pm 0.0048$
Test $0.0755\pm 0.0016$ $0.09684\pm 0.00306$ $0.12788\pm 0.0017$ $\mathbf{0.0547\pm 0.0021}$ $2.7723\pm 0.0365$ $0.1239\pm 0.0086$
@@ 6 Ablation Study
In this section, we conduct ablation studies to better understand the contribution of core components in our model architecture. Specifically, we investigate the effect of replacing the final layer with a fully connected layer (Section 6.1), substituting Depthwise convolution layers in the BinConv block with standard convolution layers (Section 6.2), and using DyTanh instead of layer normalization (Section 6.3). Each component is modified independently, while the rest of the architecture and hyperparameters remain unchanged from the standard BinConv model. During the prediction stage, we used probabilistic forecasting to measure the difference between both probabilistic and point estimates. To save computation time, we ran all modified models using a single seed and compared their performance with the average performance of BinConv.
@@
Table 7: Comparison between the standard BinConv model and its architectural variations. The notation w/ indicates that the standard BinConv is modified by replacing a specific component: w/ FC Layer uses a fully connected layer instead of the final convolution, w/ Standard Conv replaces Depthwise convolutions with standard ones, and w/ LayerNorm substitutes DyTanh with layer normalization.
Dataset BinConv w/ FC Layer w/ Standard Conv w/ LayerNorm
CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
m4_daily $0.0327$ $0.0382$ $0.0411$ $0.0468$ $0.0292$ $0.0343$ $0.0311$ $0.0365$
m4_weekly $0.0902$ $0.0972$ $0.0870$ $0.0943$ $0.1028$ $0.1112$ $0.0847$ $0.0921$
tourism_monthly $0.1824$ $0.1955$ $0.1756$ $0.1910$ $0.1801$ $0.1945$ $0.1761$ $0.1889$
@@
Dataset BinConv w/ FC Layer w/ Standard Conv w/ LayerNorm
m4_daily $20.173$ $1019.540$ $51.679$ $20.170$
m4_weekly $17.680$ $1017.164$ $44.830$ $17.677$
tourism_monthly $54.013$ $1052.210$ $146.899$ $54.010$
6.1 Using Linear Layer in the Output
When designing the BinConv architecture, we deliberately avoided using fully connected layers. There are two main reasons for this. First, fully connected layers introduce significantly more trainable parameters compared to convolutional layers, especially considering that the our model’s output lies in a high-dimensional space. Second, because of the design of BCE and the way we compute the loss during training, using a fully connected layer, especially when it serves as the final layer, may limit BinConv’s ability to forecast values on the test set that fall outside the range observed during training. We elaborate on this in detail in Appendix B.
Nevertheless, since fully connected layers are a popular design choice, we evaluate the performance of BinConv when a fully connected layer is used as the final layer. To implement this variant, we replace the final 1D convolutional layer as follows. The output of the BinConv blocks is a vector $h\in\mathbb{R}^{C\times D}$, where $C$ is the context length and $D$ is the number of bins. We first average $h$ over the context dimension, resulting in a vector $\bar{h}\in\mathbb{R}^{D}$. A fully connected layer is then applied to map $\mathbb{R}^{D}$ to $\mathbb{R}^{D}$.
In Table 7, we report the evaluation metrics for this variant in the column labeled “w/ FC Layer”. As we expected, the performance degrades significantly. Furthermore, as shown in Table 8, adding a fully connected layer dramatically increases the number of trainable parameters, making the model less efficient.
6.2 Using Depthwise instead of Standard Convolution
@@
The performance of BinConv with standard 1D convolutional layers in place of depthwise layers is reported in Table 7, in the column labeled “w/ Standard Conv”. Replacing depthwise convolutions with standard ones does not improve performance, but it does reduce model efficiency, as shown in Table 8.
6.3 Using Layer Normalization instead of DyTanh
In [zhu2025transformers], the authors proposed using DyTanh instead of layer normalization for transformer models. They demonstrated empirically that it yields comparable or better results in computer vision (CV) and natural language processing (NLP) tasks.
Inspired by this finding, we replace layer normalization with DyTanh in BinConv. In this section, we assess whether substituting DyTanh with traditional layer normalization followed by a ReLU activation affects performance. As shown in Table 7, the performance of this model variant is similar to or better than that of the standard BinConv, indicating that layer normalization is also a valid choice for our architecture. The number of parameters is also comparable (Table 8); DyTanh introduces only one additional parameter compared to the normalization layer, which is a negligible difference.
@@ 7 Conclusion
In this paper, we proposed using Binary Cumulative Encoding (BCE) as a discrete quantization technique for time series, and we introduced BinConv, a convolutional architecture specifically designed for efficient processing of BCE vectors. Extensive evaluation on the publicly available ProbTS benchmark demonstrates that our method achieves state-of-the-art performance across diverse domains.
Despite these promising results, our method has several limitations. First, BCE discretizes time series values, which requires approximating the precision of continuous data and can result in information loss. Second, selecting the number of bins introduces a trade-off: too few bins reduce representational fidelity, while too many increase computational complexity. Third, the current design is limited to univariate time series, which necessitates modeling each variable independently and prevents capturing cross-series interaction.
@@
[…]
Appendix A Evaluation on Additional Datasets
In this section, we report results for long-term forecasting on several datasets commonly used in the research community: ETTh1, ETTh2, ETTm1, ETTm2, and Illness [zhang2024probts]. For baseline models, we used the optimal set of hyperparameters from [zhang2024probts], while for BinConv we used the default configuration without any fine-tuning (see Table 10). The context length and prediction length were set to $36$ for the Illness dataset and $96$ for the others.
For each dataset and model, we ran only a single seed due to the computational cost of evaluation—particularly for models like CSDI and TimeGrad, which can require up to 20 hours for training and evaluation. All datasets are multivariate and contain 7 variables each. Since BinConv currently supports only univariate forecasting, and extending it to multivariate settings is left for future work, we forecasted each time series independently using the sampling procedure described in Section 4.2. Table 11 reports the total number of parameters for each model. BinConv is less parameter-efficient in the multivariate setting than in the univariate case, as it maintains a separate model for each time series in the dataset.
In Table 9, we report CRPS and NMAE for all baseline models and BinConv. As there is no clear winner across all datasets, we additionally report model rankings for both metrics in Table 12. Rankings were calculated based on relative performance: the best-performing model on each dataset received rank 1, the second-best received rank 2, and so on.
Despite the fact that BinConv’s hyperparameters were not tuned and the model is not specifically designed for multivariate forecasting, it achieved the best average rank based on the CRPS metric and the second-best rank based on the NMAE metric.
Table 9: Forecasting performance (CRPS and NMAE) across different models and datasets. All datasets use context length 96 and prediction length 96 (except Illness, it uses 36 for both).
Dataset DLinear PatchTST GRU-NVP TimeGrad CSDI BinConv
CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
ETTh1 0.3339 0.3339 0.3218 0.3218 0.3435 0.4383 0.2930 0.3867 0.3830 0.4871 0.2969 0.3201
ETTh2 0.2319 0.2319 0.1763 0.1763 0.2623 0.3100 0.2059 0.2574 0.2291 0.3004 0.1709 0.1853
ETTm1 0.2663 0.2663 0.2681 0.2681 0.4029 0.5018 0.4201 0.5153 0.2397 0.3124 0.2922 0.3141
ETTm2 0.1431 0.1431 0.1337 0.1337 0.2106 0.2618 0.1595 0.1930 0.1116 0.1413 0.1329 0.1456
Illness 0.1869 0.1869 0.1103 0.1103 0.0750 0.0828 0.0643 0.0868 0.3077 0.3237 0.1187 0.1447
Table 10: Default hyperparameters used for BinConv across long term datasets. $C$ denotes the context length.
Hyperparameter Value
Number of bins ($D$) $500$
Minimum bin value ($b_{0}$) $-5$
Maximum bin value ($b_{D}$) $5$
Number of output channels ($K$) $C$
Kernel size of 2D convolution in BinConv block ($s_{1}$) $3$
Kernel size of 1D convolution in BinConv block ($s_{2}$) $3$
Kernel size of final 1D convolution ($s_{3}$) $51$
Number of BinConv blocks ($M$) $3$
Dropout $0.35$
Learning rate $0.001$
Table 11: Number of trainable parameters (in thousands) for each model across datasets.
Dataset DLinear PatchTST GRU-NVP TimeGrad CSDI BinConv
ETTh1 130.368 148.240 209.059 302.071 430.001 637.663
ETTh2 18.624 33.616 414.983 115.815 430.001 637.663
ETTm1 130.368 1347.872 264.263 297.079 430.001 637.663
ETTm2 18.624 1347.872 412.459 70.199 430.001 637.663
Illness 2.664 57.836 588.039 69.623 430.001 103.063
Table 12: Model rankings by NMAE and CRPS (lower is better). Rank 1 = best performance. Averages are computed across all datasets.
Dataset DLinear PatchTST GRU-NVP TimeGrad CSDI BinConv
NMAE Rank
ETTh1 3 2 5 4 6 1
ETTh2 3 1 6 4 5 2
ETTm1 1 2 5 6 3 4
ETTm2 3 1 6 5 2 4
Illness 5 3 1 2 6 4
Average 3.0 1.8 4.8 4.2 4.4 3.0
CRPS Rank
ETTh1 4 3 5 1 6 2
ETTh2 5 2 6 3 4 1
ETTm1 2 3 5 6 1 4
ETTm2 4 3 6 5 1 2
Illness 5 3 2 1 6 4
Average 4.0 2.8 4.8 3.2 3.6 2.6
Appendix B Qualitative Analysis
B.1 The Ability to Capture the Trend
In Section 6.1, we mentioned that replacing the convolutional layer with a fully connected layer as the final layer results in the model’s inability to forecast values outside the training range. This occurs because, unlike convolutional layers, fully connected layers do not share weights. As a result, each component of the output vector is associated with its own set of learnable parameters that influence only that component.
A direct consequence of this is that if the last $K$ components of the labels during training are always zero, the network can trivially learn to predict zeros for those components, thereby failing to extrapolate any upward or downward trends.
To illustrate this, we generated a synthetic dataset with a linear trend:
$s_{t}=(100+1.5\cdot t)(1+\sigma_{t}),\quad t=0,1,\dots,T-1,$
where $\sigma_{t}\sim\mathcal{N}(0,10^{-4})$. We set $T=144$, using the first 120 values for training and the remaining 24 for testing. We then trained both the original BinConv and a variant with a fully connected output layer (BinConv w/ FC), and forecasted the final 48 values. The hyperparameters used for BinConv are provided in Table 13. Context length was set to $72$, and the prediction horizon to $24$. The BinConv w/ FC variant was constructed as described in Section 6.1. For simplicity, during forecasting, we did not sample from the output distribution but instead used the BinConv Argmax procedure from Section 4.2.
Figure 5 shows the forecast for the last 48 values alongside the ground truth. As the test phase begins, BinConv w/ FC outputs a constant value corresponding to the maximum it observed during training, failing to capture the trend. In contrast, BinConv successfully extrapolates the linear trend and achieves accurate forecasts.
Table 13: Model Parameters for Synthetic Datasets
Parameter Value
Learning Rate (lr) $0.1$
Dropout $0.0$
Number of Bins $500$
Minimum bin value ($b_{0}$) -2
Maximum bin value ($b_{D}$) 2
Number of Blocks ($K$) $1$
Kernel Size of 2D convolution in BinConv block ($s_{1}$) $3$
Kernel Size of 1D convolution in BinConv block ($s_{2}$) $3$
Kernel Size of final 1D convolution ($s_{3}$) $51$
Number of Filters (2D) $72$
Number of Filters (1D) $72$
B.2 The Ability to Capture the Seasonality
In this section, we illustrate that BinConv is capable of capturing seasonality. To demonstrate this, we generate a synthetic dataset over discrete time steps $t=0,1,\dots,T-1$. The data follows a sinusoidal seasonal pattern with additive Gaussian noise:
$s_{t}=100+A\cdot\sin\left(\frac{2\pi t}{P}\right)+\varepsilon_{t},\quad t=0,1,\dots,T-1,$
where $A=10$, $P=12$, $\varepsilon_{t}\sim\mathcal{N}(0,1)$, and $T=144$. The model hyperparameters are reported in Table 13. The context length is set to 72 and the prediction horizon to 24. The model was trained on the first 120 data values and used to forecast the last 48. The first 24 test values were seen during training, while the final 24 values were not.
Figure 6 presents the forecast results, which demonstrate that BinConv successfully captures the seasonal structure of the data.
Figure 6: Comparison of forecasting methods against observed values. The vertical line indicates the start of the test data.
Appendix C Evaluation Metrics
In our experiments, we employ the normalized mean absolute error (NMAE) for evaluating point forecasts, and the continuous ranked probability score (CRPS) for evaluating probabilistic forecasts. These metrics are widely adopted in time series forecasting research [zhang2024probts, chen2020probabilistic, ansari2024chronos, rasul2021multivariate].
NMAE is defined as:
where $K$ denotes the dimensionality of the multivariate time series (and is $1$ in the univariate case), $T$ is the prediction length (i.e., forecasting horizon), $x_{t}^{k}$ is the observed target value, and $\hat{x}_{t}^{k}$ is the predicted value.
CRPS is given by:
[…]
where $\Lambda_{\alpha}(q,z)$ is the quantile loss function at quantile level $\alpha$, defined as:
$\Lambda_{\alpha}(q,z):=\left(\alpha-\mathbb{I}{\{z<q\}}\right)(z-q),$
and $\mathbb{I}{\{z<q\}}$ is the indicator function, which equals $1$ if $z<q$ and 00 otherwise.
To approximate CRPS for discrete data, we use 19 quantile levels ranging from $0.05$ to $0.95$ in steps of $0.05$, leading to:
$\operatorname{CRPS}(F^{-1},x)\approx\sum_{i=1}^{19}2\Lambda_{i\times 0.05}\left(F^{-1}(i\times 0.05),x\right).$
To estimate the empirical quantile function $F^{-1}$, we sample 100 forecasts. For probabilistic forecasting models, the point forecast is computed as the average of these sampled forecasts. In contrast, for point forecasting models, all quantiles are presented by the single predicted value. In this case, it can be shown that the NMAE is equivalent to the CRPS, leading to identical values for both metrics.

### v3

@@ 5.3 Results
Univariate Datasets
Table 2: Default hyperparameters used for BinConv. $C$ denotes the context length. The only difference between univariate and multivariate setups is the number of bins.
Hyperparameter Univariate Multivariate
Number of bins ($D$) $1000$ $500$
Minimum bin value ($b_{0}$) 5-5 5-5
Maximum bin value ($b_{D}$) 55 55
Number of channels ($K$) $C$ $C$
K-size, conv-2D ($s_{1}$) 33 33
K-size, conv-1D ($s_{2}$) 33 33
K-size, final conv-1D ($s_{3}$) $51$ $51$
Number of blocks ($M$) 33 33
Dropout $0.35$ $0.35$
Learning rate $0.001$ $0.001$
The comparison between models is presented in Table 3. We trained and evaluated all models using five different random seeds, and report the minimum, average, and maximum scores, as well as the standard deviation for each model. The results are summarized as follows:
  • For every dataset and baseline, the average score is worse than the BinConv average score plus one standard deviation, indicating consistently better performance by BinConv.
  • Except for the weekly dataset where TimeGrad achieves the best minimum score, BinConv outperforms all baselines in terms of minimum, average, and maximum scores across all datasets.
  • For the CRPS metric, the improvement achieved by BinConv is statistically significant (p-value for T-test <0.05<0.05) for all datasets and baselines, except TimeGrad on the weekly dataset.
  • For the NMAE metric, the improvement is statistically significant (p-value for T-test <0.05<0.05) for all datasets and baselines, except TimeGrad on the weekly dataset and DLinear on the daily dataset.
Additionally, Table 4 compares the performance of BinConv when using argmax-based versus sampling-based forecasting (see Section 4.2 for details). The results show that argmax-based forecasting does not degrade point forecasting performance.
Table 3: CRPS and NMAE results for all models. Each metric is reported as the average (avg), minimum (min), maximum (max), and standard deviation (std) over 5 different seeds. An asterisk () over the BinConv average score indicates that BinConv statistically significantly outperforms all baselines for that metric and dataset ($p<0.05$).
Dataset Stat DLinear PatchTST GRU-NVP TimeGrad BinConv
@@
daily avg 0.0440 0.0440 0.0438 0.0438 0.0714 0.0782 0.0587 0.0696 0.0327 0.0382
weekly avg 0.1046 0.1046 0.1039 0.1039 0.0972 0.1062 0.1046 0.1283 0.0902 0.0972
tourism avg 0.2745 0.2745 0.2633 0.2633 0.3723 0.4536 0.2295 0.2968 0.1824 0.1955
daily min 0.0394 0.0394 0.0404 0.0404 0.0570 0.0649 0.0402 0.0475 0.0291 0.0341
weekly min 0.1007 0.1007 0.1020 0.1020 0.0916 0.1034 0.0800 0.0985 0.0851 0.0921
tourism min 0.2595 0.2595 0.2281 0.2281 0.2880 0.3849 0.1993 0.2596 0.1777 0.1911
daily max 0.0506 0.0506 0.0476 0.0476 0.1196 0.1259 0.0933 0.1077 0.0366 0.0427
weekly max 0.1072 0.1072 0.1078 0.1078 0.1006 0.1085 0.1718 0.2121 0.0981 0.1055
tourism max 0.3077 0.3077 0.2888 0.2888 0.5799 0.5399 0.2500 0.3186 0.1892 0.2012
daily std 0.0043 0.0043 0.0031 0.0031 0.0270 0.0266 0.0221 0.0252 0.0037 0.0043
weekly std 0.0025 0.0025 0.0023 0.0023 0.0038 0.0026 0.0389 0.0488 0.0050 0.0052
tourism std 0.0203 0.0203 0.0227 0.0227 0.1216 0.0696 0.0200 0.0235 0.0054 0.0049
Table 4: Comparison of BinConv performance using sampling versus argmax forecasting. Each metric is reported as the average (avg), minimum (min), maximum (max), and standard deviation (std) over 5 different seeds.
@@
daily avg 0.037 0.037 0.033 0.038
weekly avg 0.098 0.098 0.090 0.098
tourism avg 0.196 0.196 0.182 0.196
daily min 0.035 0.035 0.029 0.034
weekly min 0.093 0.093 0.085 0.092
tourism min 0.192 0.192 0.178 0.191
daily max 0.039 0.039 0.037 0.044
weekly max 0.105 0.105 0.098 0.106
tourism max 0.201 0.201 0.189 0.201
daily std 0.002 0.005 0.004 0.004
weekly std 0.005 0.005 0.005 0.005
tourism std 0.005 0.005 0.005 0.005
Multivariate Datasets
Table 5: Forecasting performance (CRPS and NMAE) across different models and datasets. All datasets use context length 96 and prediction horizon 96 (except Illness, which uses 36 for both). Best scores per row are in bold. Bottom row shows average rank (lower is better).
Dataset DLinear PatchTST GRU-NVP TimeGrad BinConv
CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
ETTh1 0.3339 0.3339 0.3218 0.3218 0.3435 0.4383 0.2930 0.3867 0.2969 0.3201
ETTh2 0.2319 0.2319 0.1763 0.1763 0.2623 0.3100 0.2059 0.2574 0.1709 0.1853
ETTm1 0.2663 0.2663 0.2681 0.2681 0.4029 0.5018 0.4201 0.5153 0.2922 0.3141
ETTm2 0.1431 0.1431 0.1337 0.1337 0.2106 0.2618 0.1595 0.1930 0.1329 0.1456
Illness 0.1869 0.1869 0.1103 0.1103 0.0750 0.0828 0.0643 0.0868 0.1187 0.1447
Avg. Rank 3.4 2.8 2.4 1.8 4.2 4.0 2.8 3.8 2.2 2.6
Due to the high computational cost, particularly for the TimeGrad model, which requires up to 20 hours for training and evaluation, we ran a single seed per model and multivariate dataset. Each dataset contains 7 variables. Since BinConv currently supports only univariate forecasting, and extending it to the multivariate setting is left for future work, we forecasted each time series independently using the sampling procedure described in Section 4.2.
In Table 5, we report CRPS and NMAE for all baseline models and BinConv. Since there is no consistent winner across all datasets, we additionally report average model rankings for both metrics. Rankings were computed based on relative performance: the best-performing model on each dataset was assigned rank 1, the second-best rank 2, and so on.
Remarkably, despite using untuned hyperparameters and not being explicitly designed for multivariate forecasting, BinConv achieved the best average rank based on the CRPS metric and the second-best rank based on the NMAE metric.
@@ 5.4 Efficiency of the Model
daily 1.204 20.173 47.470 107.434 141.108
weekly 1.040 17.680 43.725 107.434 141.108
tourism 3.504 54.013 155.288 107.434 141.108
@@
Train $\mathbf{0.0123\pm 0.0026}$ $0284\pm 0.0043 0.0$$284\pm 0.0043 0.0$$381\pm 0.0049 0.0$$450\pm 0.0063 0.0$$0362\pm 0.0048$
[…]
@@ 6 Ablation Study
6.1 Fully Connected Layers with BinConv
When designing the BinConv architecture, we intentionally excluded fully connected (FC) layers to maintain both strong generalization and parameter efficiency. Compared to convolutional layers, FC layers introduce significantly more trainable parameters, particularly problematic in high-dimensional output spaces. More critically, due to the design of CBE and the associated loss computation, an FC layer can restrict the model’s ability to forecast values beyond those observed during training. This limitation arises because FC layers lack weight sharing, meaning each output component is governed by an independent set of parameters. Consequently, if the last $K$ elements of a CBE target vector are frequently zero in the training set (a plausible scenario given that most targets are well below the maximum bin value $b_{D}$), the model may trivially learn to output zeros for these components. As a result, such an architecture fails to extrapolate to higher values during testing.
To illustrate this, we implemented a BinConv variant in which the final 1D convolution was replaced by an FC layer. The output of the BinConv blocks, $h\in\mathbb{R}^{C\times D}$, where $C$ is the context length and $D$ the number of bins, was averaged over the context dimension to yield $\bar{h}\in\mathbb{R}^{D}$, which was then mapped to $\mathbb{R}^{D}$ via the FC layer. Using a synthetic dataset with a linear trend,
$s_{t}=(100+1.5\,t)(1+\sigma_{t}),\quad\sigma_{t}\sim\mathcal{N}(0,10^{-4}),$
We compared forecasts from the original BinConv and its FC-layer variant. While the original model successfully extrapolated the trend, the FC variant saturated at the maximum value seen during training, failing to capture future growth (Figure 4). To highlight this issue, we used per-dataset mean scaling on a synthetic example. Per-sample mean scaling partially mitigates the problem, as higher test-time values shift the mean, enabling some extrapolation.
Table 8 reports evaluation metrics for the FC variant with per-sample scaling (column “w/ FC Layer”). Despite reasonable results, the model cannot produce CBE vectors with more active bins than seen during training, as many neurons in the final layer become inactive and output constant values. In addition, as shown in Table 9, the FC layer significantly increases the number of parameters. Overall, the FC variant remains ill-suited for real-world forecasting tasks.
6.2 Transformer with CBE
An alternative to modifying only the final layer of BinConv is to replace the entire architecture with a transformer-based model, which relies exclusively on FC layers. However, our experiments show that this substitution results in a substantial performance drop. We evaluated both encoder–decoder and decoder-only transformer variants, and in all cases the results were unsatisfactory. For example, on the weekly dataset the NMAE increased by nearly an order of magnitude compared to BinConv, reaching values as high as $0.8$ in the best-performing transformer configuration.
We attribute this degradation to the fundamental design of transformer architectures, which employ distinct weight matrices for processing each input component and generating each output component. While such flexibility is appropriate for one-hot encodings, it is ill-suited for the continuous bin encoding used in BinConv. As discussed in Section 6.1, CBE representations rarely contain vectors in which most components are exactly one or zero. This structural property favors architectures with weight sharing across components, ensuring that the model’s behavior is invariant to component position. Convolutional architectures naturally satisfy these constraints, making them more appropriate than transformers for processing CBE representations.
6.3 One-Hot Encoding with BinConv
In this section, we show that BinConv woorks poorly with one-hot encoding. We modify BinConv in a following way: we replace CBE with one-hot encoding both for input and output; replace final activation fucntion from sigmoid to softmax and utilze standard multi-class cross entropy loss. For this variation NMAE on the weekly dataset increases from 0.09 to 1.18, which is notably worse than all other models and baselines considered in this study. The explanation of that is that BinConv architecture, as most convolution architectures, relies heavily on neighbours component, when it forecast the CBE is replaced with one-hot encoding, the convolutional network struggles to extract meaningful features because most of the receptive field is filled with zeros. This severely limits the network’s learning capacity, resulting in significantly degraded performance. For example, NMAE on the weekly dataset increases from 0.09 to 1.18, which is notably worse than all other models and baselines considered in this study.
Unlike fully connected architectures such as transformers, BinConv operates over small local windows and therefore relies heavily on the relationships between neighboring components in the input encoding. This locality makes convolution particularly effective for structured encodings like CBE, where adjacent components often carry correlated information.
In contrast, for one-hot encodings, where exactly one component is equal to one and all others are zero, there is no meaningful local structure to exploit. The neighbors of the component with value one are always zero by design, making the surrounding values uninformative. In CBE, however, if a given component is equal to one, its neighboring components may still be either one or zero. For example, components to the right of a one-valued component may also be one, and components to the left of a zero-valued component may still be one. This structure enables convolutional filters to leverage informative local patterns, which is not possible with one-hot encodings.
@@ 6.3 One-Hot Encoding with BinConv
Table 8: Comparison of architectural variants against the standard BinConv in terms of NMAE. The notation w/ indicates that the standard BinConv is modified by replacing a specific component: w/FC uses a fully connected layer instead of the final convolution, w/SC replaces Depthwise convolutions with standard ones, colored arrows indicate whether a variant performs better () or worse () than BinConv.
Dataset BinConv w/FC w/SC
daily 0.0382 0.0468  0.0343
weekly 0.0972 0.0943  0.1112
tourism 0.1955 0.1910  0.1945
@@
Dataset BinConv w/FC w/SC
daily 20.173 1019.540 51.679
weekly 17.680 1017.164 44.830
tourism 54.013 1052.210 146.899
6.4 Depthwise Convolution instead of Standard Convolution
@@
The performance of BinConv with standard 1D convolutional layers in place of depthwise layers is reported in Table 8, under the column labeled “w/SC”. Replacing depthwise convolutions with standard ones does not improve performance on any of the datasets, and it reduces model efficiency, as shown in Table 9.
@@ 7 Conclusion
In this paper, we proposed using Cumulative Binary Encoding (CBE) as a discrete quantization technique for time series, and we introduced BinConv, a convolutional architecture specifically designed for efficient processing of CBE vectors. Extensive evaluation on the publicly available ProbTS benchmark demonstrates that our method achieves state-of-the-art performance on univariate time series datasets and performs no worse on multivariate ones compared to strong baseline models. Through a comprehensive ablation study, we show that neither CBE with fully connected architectures nor BinConv with one-hot encoding performs nearly as well as the combination of CBE and BinConv.
Despite these promising results, our method has several limitations. First, CBE discretizes time series values, which requires approximating the precision of continuous data and can result in information loss. Second, selecting the number of bins introduces a trade-off: too few bins reduce representational fidelity, while too many increase computational complexity. Third, the current design is limited to univariate time series, which necessitates modeling each variable independently and prevents capturing cross-series interaction.